# Looking Beyond Visible Cues: Implicit Video Question Answering via Dual-Clue Reasoning

### v1

@@ 5.3 Ablation Study
Experiments of the Dual Clues and Architecture. To demonstrate the effectiveness of incorporating context clues, we compare: (1) No guidance. (2) Utilizing Grounding-Dino groundingdino for object tracking. (3) Employing PySceneDetect PySceneDetect for scene detection, followed by SceneGraphParser SceneGraphParser to convert the context scenes into a triplet-based graph. As shown in Tab. 3, context dual clues lead to improvements of 3.20%, 2.53%, and 2.36%, respectively. This also proves that the SOTA performance of IRM is not simply achieved through video enhancement methods like partial masking videomae .
@@
55.14
0.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.9111148484848515151515454545457575757Percent (Unit 1 for 3.7k Data / 32 Frames / 5 Iters)Accuracy (%)
@@
Table 4: Visual Encoder and LLM Ablations. The greater impact of encoder indicates the importance of visual clues extraction.
     Visual Encoder      LLM      Acc
     EVA-CLIP-g eva      Vicuna-7B v1.5 vicuna      51.69
     EVA-CLIP-g eva      Mistral-7B  mistral      51.98
     UMT-L umt      Vicuna-7B v1.5 vicuna      54.31
     UMT-L umt      Mistral-7B mistral      55.14
@@
Table 5: Downstream Task Results of the Traffic-VQA Dataset. IRM’s SOTA performance in zero-shot indicates generalizability.
@@
           GPT-4o gpt4            45.83
           OpenAI-o3 o3            47.15
           PLLaVA pllava            37.45
           VideoChat2 videochat2            43.76
           DeepSeek-R1 deepseek            45.81
           NViLA nvila            46.12
           Self-IRM            46.22
           IRM            47.94 (+4.18)
Experiments of Data, Reasoning Time and Frames. As shown in Fig. 11, accuracy continues to rise while increasing the data volume during SFT, similar to the post-training scaling law aligner ; scaling .
We also study the relationship between reasoning time and accuracy during inference. By iteratively enhancing the visual and textual clues, as the green line in Fig. 11, performance improves with the additional 1-3 iterations (from 55.14 to 55.59, 55.73, 56.04, indicating the inference scaling law with reasoning time time ; scaling . Specifically, increasing iterations results in a better performance compared to traditional reasoning methods such as Greedy Search greedy of 55.65 and Majority Voting vote of 55.87 while maintaining the same FLOPs. Moreover, IRM reaches optimal performance with 8 frames considering efficiency, which may be related to the pre-training process.
Experiments of Visual Encoder and LLM. We default to utilizing UMT-L and Mistral-7B as our visual encoder and LLM, respectively, while also evaluating other models such as EVA-CLIP eva for the encoder and Vicuna vicuna for the LLM. As shown in Tab. 5.3, the clues extraction ability of the visual encoder has a greater impact. This shows that, in implicit reasoning, the context clues, which serve as additional textual information and enable the extraction of visual clues, are more important than the subsequent reasoning.
@@ 5.4 Generalization Evaluation
Moreover, we introduce the traffic-VQA for generalization evaluation, where models are required to trace the causes of traffic accidents. Based on the SUTD-Traffic dataset sutd , we select the “Event Forecasting” subset because it evaluates on predicting implicit outcomes. Finally, we collect 2,195 examples for generalization evaluation in a zero-shot manner. As shown in Tab. 5, IRM also performs SOTA, surpassing VideoChat2 and GPT-4o by 4.18% and 2.11%, visualization is shown in Fig. 12.
To further demonstrate the generalizability in explicit questions, we evaluate IRM on four sub-tasks of the widely accepted video understanding benchmark MVbench videochat2 (detailed in the Appendix.Sec. G). The results show that IRM outperforms VideoChat2 by 1.88% on accuracy. In particular, IRM exhibits a significant performance gain in “Action Prediction” (+8.6%) and “Moving Direction” (+4.1%), indicating that implicit video reasoning is inherently connected to predicting unknown future events.
@@ 6 Conclusion
We proposed a novel task, Implicit Video Question Answering (I-VQA), which focuses on video reasoning when explicit visual evidence is not directly accessible or locatable. Besides, we built the I-VQA dataset and proposed the first implicit VideoQA framework IRM, which introduces dual context clues as auxiliary reasoning paths. Experimental results demonstrate that IRM outperforms previous SOTA MLLMs, including GPT-4o and VideoChat2. Furthermore, IRM shows strong performance on advertisement understanding and traffic-VQA, highlighting its generalizability across a wide range of video reasoning tasks and its potential for video event prediction.
Appendix
Appendix A More Visualizations
@@ Appendix B Inference Speed
[…]
@@
VideoChat2 [46] 1.68 7B
PLLaVA [41] 1.79 7B
VideoLLaVA [39] 2.02 7B
Qwen2.5-VL [40] 2.27 7B
@@
Minigpt4-video [55] 8.94 7B
DeepSeek-R1 [50] ${}~{}\text{9.71}^{*}$ 671B
OpenAI-o3 [51] ${}~{}\text{10.70}^{*}$
[…]
Appendix C More Implementation Details
@@
Fine-tuning Details of Comparative Methods. For the fine-tuning of Qwen2.5-VL [40], we followed the official default implementation by freezing the visual encoder to preserve prior knowledge while keeping the LLM parameters learnable. Similarly, for VideoChat2 [46] and PLLaVA [41] fine-tuning, we also followed the official implementations: freezing the visual encoder while training the (Q-Former for VideoChat2) vision-language projection layers, and applying LoRA adapters tuning to the LLM components.
@@
Appendix D Details of Prompt
For the implicit question evaluations on the I-VQA (multi-choice) and the SUTD-Traffic [73] dataset, we design the guiding prompt for IRM and the other comparison models (e.g., GPT-4o) as follows:
"The question involves implicit visual information, with key visual evidencebeing invisible, requiring the deduction of the answer based on contextualvisual information and provided intention, action clues""Context clues are as follows: ""Clues1: XXX to XXX""Clues2: XXX by XXX""Clues3: XXX before XXX""Based on the clues, select the option that accurately addresses the question.""Question: XXX""(A) XXX ... (E) XXX""Only give the best option."
@@ Appendix D Details of Prompt
"The question involves implicit visual information, with key visual evidencebeing invisible, requiring the deduction of the answer based on contextualvisual information and provided intention, action clues""Context clues are as follows: ""Clues1: XXX to XXX""Clues2: XXX by XXX""Clues3: XXX before XXX""Question: XXX"
@@
"Please choose one advertising strategy that aligns with the intent of the adv-ertisement:(A) Social Identity(B) Concreteness(C) Anchoring and Comparison(D) Overcoming Reactance(E) Reciprocity(F) Foot-in-the-Door(G) Authority(H) Social Impact(I) Anthropomorphism(J) Scarcity(K) Social Proof(L) Unclear""Context clues are as follows: ""Clues1: XXX to XXX""Clues2: XXX by XXX""Clues3: XXX before XXX""Based on the clues, select the option that accurately addresses the question.""Only give the best option."
@@
[…]
The prompt templates we used for IRM, GPT-4o, OpenAI-o3 and other models are shown above. For more detailed code implementation, please refer to the open-source code repository https://anonymous.4open.science/r/Implicit-VQA/.
@@ Appendix E Data Clean Example
Appendix F Reasoning Examples of DeepSeek-R1
@@
DeepSeek-R1 [50] incorrectly assumes that the presence of a father’s care would prevent falling, instead choosing the option of dancing on the floor, which seems superficially related to the “floor”.
@@
Appendix G MVbench Subset Evaluation
To demonstrate that IRM still performs well in reasoning for explicit questions, we selected the following tasks from the MVbench [46]:
1) “Action Prediction” (AP) in the Action Reasoning.
2) “Moving Direction” (MD) in the Location Reasoning.
3) “Episodic Reasoning” (ER) in the Cognitive Reasoning.
4) “Object Existence” (OE) in the Object Reasoning.
As shown in Tab. 7, IRM significantly improves action prediction and motion direction judgment sub-tasks. This result is consistent with the previous findings in the evaluation of SUTD-VQA [73] dataset, suggesting that as the model’s understanding of implicit questions improves, its prediction and dynamic judgment capabilities also experience substantial enhancement. Moreover, the overall average accuracy increases by 1.88, demonstrating the strong robustness of our IRM. However, there is a slight decline in episodic reasoning and object existence inference, indicating that in tasks where visual information is particularly explicit, our model may experience some performance degradation. Future work incorporating multi-agents may mitigate these issues and bring about improvements.
Table 7: MVbench Subset Evaluation Results. IRM model brings a general improvement in tasks such as action evolution and behavior prediction in video understanding. Overall, the average accuracy increases by 1.88, demonstrating strong robustness.
Model AP MD ER OE Ave
VideoChat2 [46] 47.5 23.0 40.5 58.0 42.25
IRM 56.1 $\uparrow$ 27.1 $\uparrow$ 37.8 55.5 44.13$\uparrow$
+8.6 +4.1 -2.8 -2.5 +1.88
Appendix H Score Coherence Evaluation
@@
To ensure that the GPT-Score we use for open-ended evaluation is reasonable and consistent with traditional VideoQA, we calculated the Pearson correlation coefficient [59] between the average scores given by five annotators and the scores assigned by GPT-3.5. We found that the coefficient reached as high as 0.93 (with Pearson correlation values ranging from -1 to 1, where values closer to 1 indicate stronger correlation), demonstrating that this metric is also applicable to our proposed implicit-VQA.
Appendix I Details of the Related Datasets
I.1 Details of the GroundedVQA Datasets
@@
Next-GQA comprises 10,531 annotated evidence temporal segments, corresponding to 8,911 QA pairs and 1,557 videos from NextQA [42] which consists of a subset focused on complex temporal and causal reasoning. E.T. Bench includes 10,000 grounded QA samples derived from 7,002 self-collected videos, requiring models to demonstrate event-level comprehension and time-sensitive reasoning. REXTIME consists of 12,000 QA samples with annotated answer temporal locations, sourced from ActivityNet [43] and QV-Highlights [44]. It is specifically designed to evaluate the ability to perform temporal reasoning.
I.2 Details of the PSAV Dataset

### v2

@@ 5.3 Ablation Study
Experiments of the Dual Clues and Architecture. To demonstrate the effectiveness of incorporating context clues, we compare the performances between IRM and the following methods: (1) No guidance. (2) Utilizing Grounding-Dino [groundingdino] for object tracking. (3) Employing PySceneDetect [PySceneDetect] for scene detection, followed by SceneGraphParser [SceneGraphParser] to convert the context scenes into a triplet-based graph. As shown in Tab. 3, context dual clues lead to improvements of 3.20%, 2.53%, and 2.36%, respectively. This also proves that the SOTA performance of IRM is not simply achieved through video enhancement methods like partial masking [videomae].
@@
[…]
Table 4: Visual Encoder and LLM Ablations. The greater impact of encoder indicates the importance of visual clues.
Visual Encoder LLM Acc
EVA-CLIP-g [eva] Vicuna-7B v1.5 [vicuna] 51.69
EVA-CLIP-g [eva] Mistral-7B  [mistral] 51.98
UMT-L [umt] Vicuna-7B v1.5 [vicuna] 54.31
UMT-L [umt] Mistral-7B [mistral] 55.14
0.20.20.30.30.40.40.50.50.60.60.70.70.80.80.90.9114848515154545757Percent (Unit 1 for 3.7k Data / 32 Frames / 5 Iters)Accuracy (%)
@@
Experiments of Data, Reasoning Time and Frames. As shown in Fig. 11, accuracy continues to rise while increasing the data volume during SFT, similar to the post-training scaling law [aligner, scaling].
We also study the relationship between reasoning time and accuracy during inference. By iteratively enhancing the visual and textual clues, as the green line in Fig. 11, performance improves with the additional 1-3 iterations (from 55.14 to 55.59, 55.73, 56.04, indicating the inference scaling law with reasoning time [time, scaling]. Specifically, increasing iterations results in a better performance compared to traditional reasoning methods such as Greedy Search [greedy] of 55.65 and Majority Voting [vote] of 55.87 while maintaining the same FLOPs. Moreover, IRM reaches optimal performance with 8 frames considering efficiency, which may be related to the pre-training process.
Table 5: Robustness to Noisy Clues. The minimal performance drop of IRM during noisy situations highlights the effectiveness of the clue refinement module.
Model Vanilla Acc. Acc. with Noisy Clues Acc. Drop
VideoChat2 51.90 46.54 5.36
[…]
Table 6: Performance Breakdown by Question Type. IRM’s advantage is most pronounced on “How” questions, which require more complex reasoning.
Question Type IRM Acc GPT-4o Acc Difference
Why 57.39 57.25 -0.14
What 54.82 54.69 -0.13
[…]
Table 7: Downstream Task Results of the Traffic-VQA Dataset. In traffic-VQA with a certain degree of implicit meaning, IRM also shows strong generalization ability.
@@
   Proprietary LLM
   Claude-3.5-Sonnet [claude]    40.50
   Gemini-1.5-Pro [gemini]    44.96
   GPT-4o [gpt4]    45.83
[…]
   Open-Source LLM
   Minigpt4-video [Minigpt4-video]    30.71
   VideoLLaVA [videollava]    32.41
   PLLaVA [pllava]    37.45
[…]
   DeepSeek-R1 [deepseek]    45.81
   NViLA [nvila]    46.12
[…]
Experiments of Visual Encoder and LLM. We default to utilizing UMT-L and Mistral-7B as our visual encoder and LLM, respectively, while also evaluating other models such as EVA-CLIP [eva] for the encoder and Vicuna [vicuna] for the LLM. As shown in Tab. 5.3, the clues extraction ability of the visual encoder has a greater impact. This shows that, in implicit reasoning, the context clues, which serve as additional textual information and enable the extraction of visual clues, are more important than the subsequent reasoning.
Experiments of Noisy Situation Evaluations. To evaluate the robustness of our framework against irrelevant or distracting information, we conducted an experiment simulating scenarios with noisy contextual clues. Specifically, we augmented the set of potential clues by 50%, introducing noisy candidates randomly sourced from unrelated videos. This setup is designed to test the efficacy of our clue selection mechanism (the AIM module) in complex environments.
We compare the performance of our full IRM against the VideoChat2 baseline, which lacks a specialized clue-filtering mechanism. The results are presented in Tab. 5. The data shows that while the baseline model’s accuracy drops substantially by 5.36% in the presence of noisy clues, our IRM’s performance degrades by only 1.31%.
This quantitative experiment validates the effectiveness of our proposed AIM module in identifying and filtering out irrelevant information. The minimal performance drop demonstrates the robustness of the IRM framework in handling complex and noisy video contexts, a crucial capability for real-world applications. These results strongly support our claim that the proposed approach is robust, especially when navigating complex situations with many potential distractors.
5.4 Analysis of Performances of Different Types of Questions
To further investigate the implicit reasoning capabilities of our model, we conduct a fine-grained performance analysis by breaking down the results based on the type of question. We evaluate IRM and the strong baseline, GPT-4o, on three primary implicit question categories: “Why”, “What”, and “How”. The detailed accuracies are presented in Tab. 6.
The results indicate that while both models perform comparably on “Why” and “What” questions, with only marginal differences, IRM exhibits an obviously larger advantage on “How” questions. Specifically, IRM outperforms GPT-4o by 1.68% in this category, an advantage more than ten times greater than on the other two types. This finding is particularly salient, as “How” questions often necessitate more complex, multi-step causal reasoning. The substantial performance gap in this category provides strong quantitative evidence that IRM’s architecture offers a distinct advantage in handling intricate reasoning tasks.
@@ 5.5 Generalization Evaluation
Table 8: MVbench Subset Evaluation Results. IRM model brings a general improvement in tasks such as action evolution and behavior prediction in video understanding. Overall, the average accuracy increases by 1.88, demonstrating strong robustness in general video reasoning.
Model AP MD ER OE Ave
VideoChat2 [videochat2] 47.5 23.0 40.5 58.0 42.25
[…]
Moreover, although we have validated the excellent performance of IRM on the implicit advertisement understanding dataset, we remain curious to explore whether IRM can continue to perform better in generalized, real-world video understanding tasks involving more flexible implicit meaning. Therefore, we introduce a classical real-world video understanding dataset, Traffic-VQA, for evaluation, where models are required to trace the causes of traffic accidents. Specifically, we introduce the “Event Forecasting” subset of SUTD-Traffic dataset [sutd], since it evaluates on predicting implicit outcomes. Finally, we collect 2,195 examples for generalization evaluation in a zero-shot manner. As shown in Tab. 7, IRM also performs SOTA, surpassing VideoChat2 and GPT-4o by 4.18% and 2.11%, visualization is shown in Fig. 12. This result proves that IRM can also enhance the performance of answering real-world questions with a certain degree of implicit meaning, such as those related to prediction or planning.
Furthermore, we examine whether IRM can effectively reason over more general video questions. In these questions, the distinction between explicit and implicit types is less clear. To this end, we selected several representative tasks from the widely recognized video understanding benchmark MVbench [videochat2]:
1) “Action Prediction” (AP) in the Action Reasoning.
2) “Moving Direction” (MD) in the Location Reasoning.
3) “Episodic Reasoning” (ER) in the Cognitive Reasoning.
4) “Object Existence” (OE) in the Object Reasoning.
As shown in Tab. 8, IRM significantly improves action prediction and motion direction judgment sub-tasks. This result aligns with the findings from the evaluation on the SUTD-Traffic [sutd] dataset, indicating that as the model’s ability to handle implicit questions improves, its predictive and dynamic reasoning capabilities are also significantly enhanced. Moreover, the overall average accuracy increases by 1.88, demonstrating the strong robustness of our IRM in general videoQA tasks. However, there is a slight decline in episodic reasoning and object existence inference, indicating that in tasks where visual information is particularly explicit, our model may experience minimal performance degradation.
Overall, on highly general video understanding tasks, our IRM can still balance performance across video questions with varying degrees of implicit meaning, and it continues to provide an overall performance boost.
@@ 6 Conclusion
We introduce a novel task, Implicit Video Question Answering (I-VQA), which focuses on video reasoning in scenarios where explicit visual evidence is neither directly accessible nor easily locatable. Our findings demonstrate that current state-of-the-art Multimodal Large Language Models (MLLMs) struggle to effectively tackle implicit reasoning, underscoring the importance of advancing research in implicit video question answering. Additionally, we developed the I-VQA dataset through a robust semi-automated pipeline and proposed the first implicit VideoQA framework, IRM, which integrates dual context clues as auxiliary reasoning paths. Experimental results show that IRM outperforms leading MLLMs, such as GPT-4o, fine-tuned Qwen2.5-VL, and VideoChat2 in implicit video question answering. Moreover, IRM demonstrates strong performance across multiple VideoQA datasets—PSAV, SUTD-Traffic, and MVBench—underscoring its generalizability to real-world video question answering tasks with varying levels of implicit content.
7 Statements and Declarations
All annotations for the I-VQA dataset, including question-answer pairs and context clues annotated by GPT-4o, are now publicly available in the repository: [https://github.com/tychen-SJTU/Implicit-VideoQA].
The original video sources can be legally obtained from several benchmarks, including Next-GQA [nextgqa], E.T. Bench [etbench], and REXTIME [rextime].
This work was supported in part by the National Natural Science Foundation of China under Grant 62325109, Grant 62561160155, and Grant U21B2013, in part by the Shanghai ‘The Belt and Road’ Young Scholar Exchange under Grant 24510742000, and in part by the National Key R&D Program of China under Grant 2022ZD0160102.
Appendix A Visualizations
@@ Appendix B Inference Speed
Table 9: Inference Speed. Our IRM not only achieves SOTA performance but also operates at a relatively fast speed. * indicates with evaluation through API. $\dagger$} indicates estimation.
@@
VideoChat2 [videochat2] 1.68 7B
PLLaVA [pllava] 1.79 7B
VideoLLaVA [videollava] 2.02 7B
Qwen2.5-VL [qwen] 2.27 7B
@@
Minigpt4-video [Minigpt4-video] 8.94 7B
DeepSeek-R1 [deepseek] $\text{9.71}^{*}$ 671B
OpenAI-o3 [o3] $\text{10.70}^{*} -$
GPT-4o [gpt4o] $\text{14.11}^{*}$ $\text{200B}^{\dagger}$
Appendix C Implementation
@@
Fine-tuning Details of Comparative Methods. For the fine-tuning of Qwen2.5-VL [qwen], we followed the official default implementation by freezing the visual encoder to preserve prior knowledge while keeping the LLM parameters learnable. Similarly, for VideoChat2 [videochat2] and PLLaVA [pllava] fine-tuning, we also followed the official implementations: freezing the visual encoder while training the (Q-Former for VideoChat2) vision-language projection layers, and applying LoRA adapters tuning to the LLM components.
@@
Appendix D Prompt
For the implicit question evaluations on the I-VQA (multi-choice) and the SUTD-Traffic [sutd] dataset, we design the guiding prompt for IRM and the other comparison models (e.g., GPT-4o) as follows:
"The question involves implicit visual information, with key visual evidence being invisible, requiring the deduction of the answer based on contextual visual information and provided intention, action clues"
"Context clues are as follows: "
"Clues1: XXX to XXX"
"Clues2: XXX by XXX"
"Clues3: XXX before XXX"
"Based on the clues, select the option that accurately addresses the question."
"Question: XXX"
"(A) XXX ... (E) XXX"
"Only give the best option."
@@ Appendix D Prompt
"The question involves implicit visual information, with key visual evidence being invisible, requiring the deduction of the answer based on contextual visual information and provided intention, action clues"
"Context clues are as follows: "
"Clues1: XXX to XXX"
"Clues2: XXX by XXX"
"Clues3: XXX before XXX"
"Question: XXX"
@@
"Please choose one advertising strategy that aligns with the intent of the advertisement:
(A) Social Identity
(B) Concreteness
(C) Anchoring and Comparison
(D) Overcoming Reactance
(E) Reciprocity
(F) Foot-in-the-Door
(G) Authority
(H) Social Impact
(I) Anthropomorphism
(J) Scarcity
(K) Social Proof
(L) Unclear"
"Context clues are as follows: "
"Clues1: XXX to XXX"
"Clues2: XXX by XXX"
"Clues3: XXX before XXX"
"Based on the clues, select the option that accurately addresses the question."
"Only give the best option."
@@
[‘Human’, The question involves some implicit visual information, with key visual evidence being invisible, requiring the inference of the correct answer based on context. ’],
[‘Human’, Context clues are as follows: boy climbs onto apparatus to engage in gymnastics activity, the boy holds onto gymnastic bars to practice balance.’],
[‘Human’, Based on the clues, select the best option that accurately
addresses the question.’],
[‘Human’, Q: Please answer: What does the boy in red and black shorts do as the boy in blue shorts was hanging on the pole?
(A) walk in front of him
(B) watch the kids
(C) jumps withthem
(D) dropped on table
(E) climb on table’],
[‘Human’, Only give the best option.’]
The prompt templates we used for IRM, GPT-4o, OpenAI-o3 and other models are shown above. For more detailed code implementation, please refer to the open-source code repository https://github.com/tychen-SJTU/Implicit-VideoQA.
@@ Appendix E Data Clean Example
Appendix F Examples of DeepSeek-R1
@@
DeepSeek-R1 [deepseek] incorrectly assumes that the presence of a father’s care would prevent falling, instead choosing the option of dancing on the floor, which seems superficially related to the “floor”.
@@
Appendix G Score Coherence Evaluation
@@
To ensure that the GPT-Score we use for open-ended evaluation is reasonable and consistent with traditional VideoQA, we calculated the Pearson correlation coefficient [pearson] between the average scores given by five annotators and the scores assigned by GPT-3.5. We found that the coefficient reached as high as 0.93 (with Pearson correlation values ranging from -1 to 1, where values closer to 1 indicate stronger correlation), demonstrating that this metric is also applicable to our proposed implicit-VQA.
Appendix H Datasets
H.1 Details of the GroundedVQA Datasets
@@
Next-GQA comprises 10,531 annotated evidence temporal segments, corresponding to 8,911 QA pairs and 1,557 videos from NextQA [nextqa] which consists of a subset focused on complex temporal and causal reasoning. E.T. Bench includes 10,000 grounded QA samples derived from 7,002 self-collected videos, requiring models to demonstrate event-level comprehension and time-sensitive reasoning. REXTIME consists of 12,000 QA samples with annotated answer temporal locations, sourced from ActivityNet [acititynet] and QV-Highlights [qv-highlight]. It is specifically designed to evaluate the ability to perform temporal reasoning.
H.2 Details of the PSAV Dataset